Solve for x:
5x^2=-17x-6
I have no idea where to begin. Thanks.

Found 2 solutions by unlockmath, richwmiller:
Answer by unlockmath(1688)

Hello,
When we are dealing with equations that have x to a power in this case to the 2nd power we need to get everything to one side of the equal sign. For this one we add 17x and 6 to both sides like this:
5x^2+17x+6=-17x+17x-6+6 This turns out to be:
5x^2+17x+6=0 Now we can factor (factor simply means two numbers multiplied)
(5x+2)(x+3)=0 When you multiply these two, we get 5x^2+17x+6=0
Now we can set up this:
x+3=0 Subtract 3 from both sides so x=-3
5x+2=0 Subtract 2 and divide by 5 gives us:
x=-2/5
So we get two answers.
X=-3 and x=-2/5
Plug the answer into the original equation and it will work out.
Does this help you?
I suggest you get the book "Unlock the Mystery to Math" at www.math-unlock.com
It goes over the very basics of math in easy to read form.
RJ toftness

Answer by richwmiller(17219)

5x^2=-17x-6
we want to put it in the style of
ax^2+bx+c=0
to move everything to one side we add 17x+6 to both sides
5x^2+17x+6=0
I suggest formula the quadratic formula
x=-b+-sqrt(b^2-4ac)/(2*a)
where a=5; b=17 and c=6
so we get b^2-4ac=17^2-4*5*6
Since 17^2-4*5*6 is greater than zero we will have two solutions.
continue plugging in the values a;b and c
the two solutions are
(-17+sqrt(17^2-4*5*6))/2*5
(-17-sqrt(17^2-4*5*6))/2*5